# Neutron Star-White Dwarf Merger as One Possible Optional Source of Kilonova-like Emission: Implications for GRB 211211A

### v1

@@ 3.1 Numerical Simulations of Nucleosynthesis
Based on the method and theories of nucleosynthesis mentioned in Section 2.1, we perform numerical simulations of nucleosynthesis via SkyNet for both the disk and ejecta from WD-NS mergers. Figure 1 shows the final abundance of the ejecta for three different groups of WD-NS systems with mass groups of (0.75-1.4) $M_{\odot}$ (Model-A), (1-1.25) $M_{\odot}$ (Model-B), and (1-1.8) $M_{\odot}$ (Model-C), respectively. Moreover, we also present the final abundance of the ejecta with different temperatures (i.e., 4 GK, 5 GK, and 6 GK) for each model. For Model-A, it is found that the final abundance of ejecta for temperature $T=5$ GK is not very different from that of temperature $T=6$ GK, and the nuclides produced by $T=5$ GK are only slightly heavier than those produced by $T=6$ GK. However, the abundance of heavier elements (i.e., heavier than iron) with $T=5$ GK and $T=6$ GK is greater than that produced by $T=4$ GK. By adopting different mass distributions of WD and NS in the binary, the results of Model-B and Model-C exhibit similar trends to those observed in Model-A.
@@
Figure 1: Final abundance of the ejecta for Model-A (top panel), Model-B (middle panel), and Model-C (bottom panel) with different temperatures ($T=4$ GK, 5 GK, and 6 GK), respectively.
@@
On the other hand, at the given temperature of $T=4$ GK, we find that the abundances in all three models peak at mass number $A=56$, namely, the abundance of $\rm{}^{56}Fe$ is the highest. If this is the case, the decay chain $\rm{}^{56}Ni\rightarrow{}^{56}Co\rightarrow{}^{56}Fe$ can release a large amount of energy. In addition, the mass number of the heaviest elements it can produce is in the range of $7060<A<$ 70. At the given temperatures of $T=5$ GK and $T=6$ GK, a significant amount of elements with mass number $A=56$ are still produced, but the major difference compared with the result at $T=4$ GK is that the abundance of $\rm{}^{4}He$ is substantially higher. There are two primary ways to cause such a higher abundance of $\rm{}^{4}He$. One is that the iron peak elements can be decomposed into free nuclei (protons, neutrons) and lighter nuclei (e.g., $\alpha$-particles) when the initial temperature is higher ($T\sim 5$ GK and 6 GK). With the ejecta rapidly expanding and cooling down, the nuclear reactions, which were initially in dynamic equilibrium, are ’frozen’, and the decomposed $\alpha$-particles are retained because they are too late to recombine into heavy nuclei. The other one is that much of $\rm{}^{4}He$ can be produced by the combination of neutrons and protons at a higher temperature. With the ejecta expanding at high speed, the density of the ejecta decreases, which makes $\rm{}^{4}He$ difficult to be consumed.
@@
Moreover, by comparing the heavier elements of the three models at the same temperature, it is found that the mass number of the produced heavier nuclides follows $A_{\rm Model-A}>A_{\rm Model-C}>A_{\rm Model-B}$. This is because the entropy of these three models, which is related to the density as $s\propto{\rho}^{-1}$, is calculated using the Helmholtz equation of state (EOS). As shown in Table 1, the density of the ejecta in Model-A is the smallest among the three Models, but it has the highest entropy. Therefore, Model-A tends to produce elements with higher mass numbers compared to the other two models. Based on Figure 1, no matter which temperature or model we choose , the mass number of the heaviest produced nuclides is less than 90 (i.e., $A<90$), and a large fraction of the produced elements are around the iron peak.
@@ 3.2 Numerical Calculation of Light Curve of kilonova-like event
As mentioned in Section 3.1, the decay chain of $\rm{}^{56}Ni\rightarrow{}^{56}Co\rightarrow{}^{56}Fe$ can release a large amount of energy, which is used to heat the ejecta to power the kilonova-like emission. In order to test how bright the kilonova-like emission powered by the radioactive decay can be, following the methods presented in Section 2.3 and 2.2, we calculate the light curves of kilonova-like emission for the three models with three different temperatures, respectively. Figure 2 shows the $r-$band light curves of the kilonova-like for all three models [^1] Here, we adopt the luminosity distance $D_{\rm L}=346\rm~{}Mpc$ which is the distance of GRB 211211A (Troja et al., 2022). For the same given temperature, it is found that the kilonova-like emission of Model-B is the brightest. That is because the peak luminosity is strongly dependent on the mass of the ejecta, namely, the larger the ejecta mass, the higher the peak luminosity of the kilonova-like emission.
On the other hand, by comparing the results for different temperatures within each model, we find that the kilonova-like emission at $T=4$ GK is brighter than that at $T=5$ GK and $T=6$ GK for each model. The reason may be that the abundance of $\rm{}^{56}Ni$ at $T=4$ GK is higher than that at $T=5$ GK and $T=6$ GK during the long evolution of the nuclear reaction network, and $\rm{}^{56}Ni$ will be decomposed into $\alpha$-particles at $T=5$ GK and above, resulting in a lower abundance of $\rm{}^{56}Ni$. However, the brightness of the kilonova-like emission at $T=5$ GK is slightly higher than at $T=6$ GK (see Figure 2), indicating that the abundance of $\rm{}^{56}Ni$ does not differ significantly between these two temperatures.
@@ 4 Application to the possible kilonova emission associated with GRB 211211A
In recent years, a peculiar and nearby long-duration GRB 211211A with a redshift of $z=0.076$ has attracted great attention (Rastinejad et al., 2022; Troja et al., 2022; Yang et al., 2022; Chang et al., 2023; Gompertz et al., 2023). The lack of an associated supernova with GRB 211211A at such a low redshift, but associated with a possible kilonova emission, together with the long duration of the prompt emission, suggest that its physical origin may be a compact binary star merger, especially, a merger of a WD and an NS (Yang et al., 2022; Zhong et al., 2023). In Section 3, we briefly discussed the heavy elements nucleosynthesis and kilonova-like emission from NS-WD merger. In this Section, from the perspective of heavy elements nucleosynthesis, we will discuss whether the NS-WD merger can indeed power a confirmed observational case such as a possible kilonova emission associated with GRB 211211A.
@@
In order to compare the calculated kilonova-like emission with that of afterglow-subtracted observations of possible kilonova emission in GRB 211211A, we adopt Model-B, which is the brightest model for powering kilonova-like emission. By collecting the data of possible kilonova emission (or even upper limits) with afterglow-subtracted of GRB 211211A from Yang et al. (2022), we plot the magnitude of possible kilonova emission in r-band for GRB 211211A to compare with the calculated kilonova-like emission from Model-B (see Figure 3). The magnitude of kilonova-like emission at the temperature of $T=4$ GK reaches its peak at about 3 days, and that of $T=5$ GK and $T=6$ GK peaks at about 2.5 days. We find that the observations of the possible kilonova emission of GRB 211211A are lower than those of our calculated model after 1 day. It means that the calculated kilonova-like emission from radioactive decay of the nuclei is sufficient to provide the energy to power the possible kilonova emission associated with GRB 211211A after 1 day of the burst. However, the calculated luminosity from the theoretical model cannot reach such a high luminosity at the early stage, and the afterglow-subtracted observations of the possible kilonova emission associated with GRB 211211A are brighter. It suggests that a single energy source (only radioactive decay) could not provide enough energy to power the possible kilonova emission of GRB 211211A before 1 day after burst. In other words, it may be necessary to add other energy sources, such as the spin-down energy of a newborn magnetar (Yang et al., 2022; Zhong et al., 2023; Wang et al., 2024).
@@
If GRB 211211A is indeed originated from the merger of a WD and an NS, the possible kilonova emission following GRB 211211A would be powered not only by radioactive decay of the nuclei, but also by the spin-down energy from the newborn magnetar, or other energy sources. If this is the case, the remnant of the WD-NS merger should be a supramassive or stable magnetar. Otherwise, it is difficult to explain the early possible kilonova emission following GRB 211211A by invoking the merger of WD and NS.
@@
Since the brightness of kilonova-like emission depends on the temperature and ejecta mass, and the latter is related to the masses of the WD and the NS. The higher mass of the WD and lower mass of the NS in the WD-NS binary system can result in a larger ejecta mass. We calculate the kilonova-like emission again by adopting $T=3$ GK, and find that the brightness of the kilonova-like emission is still fainter than that of $T=4$ GK. Moreover, we also adopt the equal mass of WD and NS (e.g., $M_{\rm WD}=M_{\rm NS}=1.25\rm~{}M_{\odot}$) to do the simulation and calculate the kilonova-like emission again, and also find that the calculated kilonova-like emission is still fainter than that of observations of possible kilonova associated with GRB 211211A at the early stage.
@@
Figure 3: Comparison of calculated kilonova-like emission in r-band for Model B with three different temperatures at $\rm D_{L}=346Mpc$ and the observations of the possible kilonova associated with GRB 211211A. The data (orange points) and upper limits (arrows) of the afterglow-subtracted observations are taken from (Yang et al., 2022).
@@ 5 Conclusion and Discussion
The origin of kilonova-like emission following the long-duration GRBs is currently the subject of much discussion, and mergers of NS and WD are also proposed to naturally explain the long-duration of GRBs (Yang et al., 2022; Zhong et al., 2023; Wang et al., 2024). However, the nucleosynthesis of heavy elements and the brightness of the kilonova-like emission during the mergers of NS and WD are still in debate. In this paper, by adopting three groups of NS and WD with different masses and three different temperatures (Margalit & Metzger, 2016; Zenati et al., 2019; Kaltenborn et al., 2023), we have studied the heavy elements that are synthesized in NS-WD mergers based on the nucleosynthesis simulations, and then compared the results with the observed solid case of possible kilonova emission of GRB 211211A. We find the following interesting results.
@@
  • Kilonova-like light curves are calculated for three models at different temperatures (4 GK, 5 GK, and 6 GK), and it is found that Model-B shows the brightest emission at each given temperature due to its larger ejecta mass. By comparing with adopting different temperatures of each model, the kilonova-like emission at $T=4$ GK is brighter than that of $T=5$ GK and $T=6$ GK, which is likely due to the higher $\rm{}^{56}Ni$ production at $T=4$ GK.
@@
  • By comparing the calculated kilonova-like emission with that of afterglow-subtracted observations of possible kilonova emission in GRB 211211A, the merger of NS and WD cannot be ruled out as the origin of GRB 211211A to produce the possible kilonova emission if the remnant of the WD-NS merger is a supramassive or stable magnetar. Otherwise, it is difficult to explain the early possible kilonova emission following GRB 211211A by invoking the merger of WD and NS.
@@
We also compare our results with those of other works derived from different models. In Zenati et al. (2020), they adopted 3D hydrodynamical simulations and 2D hydrodynamical-thermonuclear simulations to model the mergers of NSs with CO-WDs, which produce faint transient events. In their results, the predicted transient events are much fainter and faster evolving than those of typical Type Ia SNe. For the peak luminosity, the luminosity at $T=4$ GK in our result is not much different from that of their work, with the unit of magnitude. However, in terms of the time of the peak, the peak time of our results is earlier than that of their work. The reasons for such a difference may be due to different heating rates, or different models for calculating the radiative transfer, or differences in expansion velocity and ejecta masses. In Morán-Fraile et al. (2024), the 3D-MHD simulations of the merger were conducted by using the magnetohydrodynamic moving mesh code AREPO. They claimed that the luminosity of the transients produced by NS-WD mergers could be larger than that of typical kilonovae ($\sim 10^{41}\rm~{}erg/s$), which is also consistent with our results from the aspect of the total luminosity.
@@
In any case, it is difficult to firmly identify the origin of the GRB 211211A-like event only based on its multi-wavelength kilonova-like emission. The best way to identify the merger system is to combine with other observational characteristics, such as direct GW detection, or more detailed spectroscopic analysis to look for spectral line features produced by heavier elements, as seen in examples like AT2017gfo, AT2018kzr, and AT2023vfi (Gillanders et al., 2020, 2023; Gillanders & Smartt, 2025; Liu et al., 2025).
@@
Moreover, GRB 230307A is another interesting event which is similar to that of GRB 211211A (Gillanders et al., 2023; Du et al., 2024; Levan et al., 2024; Yang et al., 2024; Sun et al., 2025), and the characteristic of soft X-ray emission suggest that the central engine of GRB 230307A is a magnetar from a binary NS merger (Du et al., 2024; Zhong et al., 2024; Sun et al., 2025). Wang et al. (2024) proposed that the origin of GRB 230307A is also a NS–WD merger, and the late-time kilonova-like emission is powered by the spin-down of the magnetar and the radioactive decay of $\rm{}^{56}Ni$. More interestingly, the prominent emission feature of GRB 230307A at $\sim 2.1\mu m$ detected by James Webb Space Telescope (JWST) is suggested to be likely attributed to Te III with $A=128$ (Gillanders & Smartt, 2025). Such an element is much heavier than the elements obtained by our simulations from the merger of NS and WD. If such an emission line in GRB 230307A is indeed real, it means that at least GRB 230307A originated from the merger of NS and WD is very difficult to explain the observed emission line of Te III in the spectrum. In other words, once one detects the emission line of heavier elements ($A>90$) in the spectrum of a GRB 230307A-like event in the feature, the origin of the mergers of NS and WD would be ruled out.
@@
Meanwhile, it is worth noting that the system of NS-WD merger is not the only progenitor to explain the type of LGRBs associated with possible kilonova emission. A number of other possible origins are also proposed, such as neutron star-neutron star mergers (NS-NS; Gottlieb et al. (2023)), neutron star-black hole mergers (NS-BH; Rastinejad et al. (2022)), WD-BH mergers (Fryer et al., 1999; Lloyd-Ronning et al., 2024), massive accretion disks around black hole (Gottlieb et al., 2025), and low-mass collapsar as the progenitor (Zenati et al., 2020; An et al., 2023; Barnes & Metzger, 2023; Dean & Fernández, 2024). Further multi-messenger observations will be essential to distinguish between these progenitor scenarios and to establish a more comprehensive understanding of the origin of this type of LGRBs.
Our simulation also poses some curious questions. One concern is that the mass number of the heaviest element produced by radioactive decay at $T=5$ GK is larger than that of $T=6$ GK. In general, the reaction rate of nuclides should increase with the rise in temperature. However, the increase of reaction rate for different nuclides may differ when the temperature rises by 1 GK. Thus, a possible explanation for the above question is that the increase in reaction rates for the nuclides with mass numbers in the range $6030<A<$ 60 is larger than that for those with mass numbers in the range $8060<A<$ 80 when the temperature increases from $T=5$ GK to $T=6$ GK. If this is the case, it can result in more nuclides with mass numbers in the range $6030<A<$ 60, thereby suppressing the production of heavier elements.

### v2

@@ 3.1 Numerical Simulations of Nucleosynthesis
Based on the method and theories of nucleosynthesis mentioned in Section 2.1, we perform numerical simulations of nucleosynthesis via SkyNet for both the disk and ejecta from WD-NS mergers. Figure 1 shows the final abundance of the ejecta for three different groups of WD-NS systems with mass groups of (0.75-1.4) $M_{\odot}$ (Model-A), (1-1.25) $M_{\odot}$ (Model-B), and (1-1.8) $M_{\odot}$ (Model-C), respectively. Moreover, we also present the final abundance of the ejecta with different temperatures (i.e., 4, 5, and 6 GK) for each model. For Model-A, it is found that the final abundance of ejecta for temperature $T=5$ GK is not very different from that of temperature $T=6$ GK, and the nuclides produced by $T=5$ GK are only slightly heavier than those produced by $T=6$ GK. However, the abundance of heavier elements (i.e., heavier than iron) with $T=5$ GK and $T=6$ GK is greater than that produced by $T=4$ GK. By adopting different mass distributions of WD and NS in the binary, the results of Model-B and Model-C exhibit similar trends to those observed in Model-A.
@@
Figure 1: Final abundance of the ejecta for Model-A (top panel), Model-B (middle panel), and Model-C (bottom panel) with different temperatures ($T=4$, 5, and 6 GK), respectively.
@@
On the other hand, at the given temperature of $T=4$ GK, we find that the abundances in all three models peak at mass number $A=56$, namely, the abundance of $\rm{}^{56}Fe$ is the highest. If this is the case, the decay chain $\rm{}^{56}Ni\rightarrow{}^{56}Co\rightarrow{}^{56}Fe$ can release a large amount of energy. In addition, the mass number of the heaviest elements it can produce is in the range of $60<A<70$. At the given temperatures of $T=5$ GK and $T=6$ GK, a significant amount of elements with mass number $A=56$ are still produced, but the major difference compared with the result at $T=4$ GK is that the abundance of $\rm{}^{4}He$ is substantially higher. There are two primary ways to cause such a higher abundance of $\rm{}^{4}He$. One is that the iron peak elements can be decomposed into free nuclei (protons, neutrons) and lighter nuclei (e.g., $\alpha$ particles) when the initial temperature is higher ($T\sim 5$ and 6 GK). With the ejecta rapidly expanding and cooling down, the nuclear reactions, which were initially in dynamic equilibrium, are ”frozen”, and the decomposed $\alpha$ particles are retained because they are too late to recombine into heavy nuclei. The other one is that much of $\rm{}^{4}He$ can be produced by the combination of neutrons and protons at a higher temperature. With the ejecta expanding at high speed, the density of the ejecta decreases, which makes $\rm{}^{4}He$ difficult to be consumed.
@@
Moreover, by comparing the heavier elements of the three models at the same temperature, it is found that the mass number of the produced heavier nuclides follows $A_{\rm Model-A}>A_{\rm Model-C}>A_{\rm Model-B}$. This is because the entropy of these three models, which is related to the density as $s\propto{\rho}^{-1}$, is calculated using the Helmholtz EOS. As shown in Table 1, the density of the ejecta in Model-A is the smallest among the three models, but it has the highest entropy. Therefore, Model-A tends to produce elements with higher mass numbers compared to the other two models. Based on Figure 1, no matter which temperature or model we choose , the mass number of the heaviest produced nuclides is less than 90 (i.e., $A<90$), and a large fraction of the produced elements are around the iron peak.
@@ 3.2 Numerical Calculation of Light Curve of kilonova-like event
As mentioned in Section 3.1, the decay chain of $\rm{}^{56}Ni\rightarrow{}^{56}Co\rightarrow{}^{56}Fe$ can release a large amount of energy, which is used to heat the ejecta to power the kilonova-like emission. In order to test how bright the kilonova-like emission powered by the radioactive decay can be, following the methods presented in Section 2.3 and 2.2, we calculate the light curves of kilonova-like emission for the three models with three different temperatures, respectively. Figure 2 shows the $r-$band light curves of the kilonova-like emission for all three models [^1] Here, we adopt the luminosity distance $D_{\rm L}=346\rm~Mpc$ which is the distance of GRB 211211A (Troja et al., 2022). For the same given temperature, it is found that the kilonova-like emission of Model-B is the brightest. That is because the peak luminosity is strongly dependent on the mass of the ejecta; namely, the larger the ejecta mass, the higher the peak luminosity of the kilonova-like emission.
On the other hand, by comparing the results for different temperatures within each model, we find that the kilonova-like emission at $T=4$ GK is brighter than that at $T=5$ GK and $T=6$ GK for each model. The reason may be that the abundance of $\rm{}^{56}Ni$ at $T=4$ GK is higher than that at $T=5$ GK and $T=6$ GK during the long evolution of the nuclear reaction network, and $\rm{}^{56}Ni$ will be decomposed into $\alpha$ particles at $T=5$ GK and above, resulting in a lower abundance of $\rm{}^{56}Ni$. However, the brightness of the kilonova-like emission at $T=5$ GK is slightly higher than at $T=6$ GK (see Figure 2), indicating that the abundance of $\rm{}^{56}Ni$ does not differ significantly between these two temperatures.
@@ 4 Application to the possible kilonova emission associated with GRB 211211A
In recent years, the peculiar and nearby long-duration GRB 211211A with a redshift of $z=0.076$ has attracted great attention (Rastinejad et al., 2022; Troja et al., 2022; Yang et al., 2022; Chang et al., 2023; Gompertz et al., 2023). The lack of an associated SN with GRB 211211A at such a low redshift, but associated with a possible kilonova emission, together with the long duration of the prompt emission, suggest that its physical origin may be a compact binary star merger, especially, a merger of a WD and an NS (Yang et al., 2022; Zhong et al., 2023). In Section 3, we briefly discussed the heavy elements nucleosynthesis and kilonova-like emission from NS-WD mergers. In this Section, from the perspective of heavy elements nucleosynthesis, we will discuss whether the NS-WD merger can indeed power a kilonova-like emission associated with GRB 211211A.
@@
In order to compare the calculated kilonova-like emission with that of afterglow-subtracted observations of possible kilonova emission in GRB 211211A, we adopt Model-B, which is the brightest model for powering kilonova-like emission. By collecting the data of possible kilonova emission (or even upper limits) with afterglow-subtracted observation of GRB 211211A from Yang et al. (2022), we plot the magnitude of possible kilonova emission in r-band for GRB 211211A to compare with the calculated kilonova-like emission from Model-B (see Figure 3). The magnitude of kilonova-like emission at the temperature of $T=4$ GK reaches its peak at about 3 days, and that of $T=5$ GK and $T=6$ GK peaks at about 2.5 days. We find that the observations of the possible kilonova emission of GRB 211211A are lower than those of our calculated model after 1 day. It means that the calculated kilonova-like emission from radioactive decay of the nuclei is sufficient to provide the energy to power the possible kilonova emission associated with GRB 211211A after 1 day of the burst. However, the calculated luminosity from the theoretical model cannot reach such a high luminosity at the early stage, and the afterglow-subtracted observations of the possible kilonova emission associated with GRB 211211A are brighter. It suggests that a single energy source (only radioactive decay) could not provide enough energy to power the possible kilonova emission of GRB 211211A before 1 day after burst. In other words, it may be necessary to add other energy sources, such as the spin-down energy of a newborn magnetar (Yang et al., 2022; Zhong et al., 2023; Wang et al., 2024).
@@
If GRB 211211A is indeed originated from the merger of a WD and an NS, the possible kilonova emission following GRB 211211A would be powered not only by radioactive decay of the nuclei but also by the spin-down energy from the newborn magnetar or other energy sources. If this is the case, the remnant of the WD-NS merger should be a supramassive or stable magnetar. Otherwise, it is difficult to explain the early possible kilonova emission following GRB 211211A by invoking the merger of WD and NS.
@@
Since the brightness of kilonova-like emission depends on the temperature and ejecta mass, the latter is related to the masses of the WD and the NS, the higher mass of the WD and lower mass of the NS in the WD-NS binary system can result in a larger ejecta mass. We calculate the kilonova-like emission again by adopting $T=3$ GK, and find that the brightness of the kilonova-like emission is still fainter than that of $T=4$ GK. Moreover, we also adopt the equal mass of a WD and an NS (e.g., $M_{\rm WD}=M_{\rm NS}=1.25~M_{\odot}$) to do the simulation and calculate the kilonova-like emission again and also find that the calculated kilonova-like emission is still fainter than that of observations of possible kilonova associated with GRB 211211A at the early stage.
@@
Figure 3: Comparison of calculated kilonova-like emission in r-band for Model-B with three different temperatures at $\rm D_{L}=346Mpc$ and the observations of the possible kilonova associated with GRB 211211A. The data (orange points) and upper limits (arrows) of the afterglow-subtracted observations are taken from (Yang et al., 2022).
@@ 5 Conclusion and Discussion
The origin of kilonova-like emission following the LGRBs is currently the subject of much discussion, and mergers of an NS and a WD are also proposed to naturally explain the long duration of GRBs (Yang et al., 2022; Zhong et al., 2023; Wang et al., 2024). However, the nucleosynthesis of heavy elements and the brightness of the kilonova-like emission during the mergers of NS and WD are still in debate. In this Letter, by adopting three groups of NSs and WDs with different masses and three different temperatures (Margalit & Metzger, 2016; Zenati et al., 2019; Kaltenborn et al., 2023), we studied the heavy elements that are synthesized in NS-WD mergers based on the nucleosynthesis simulations and then compared the results with the observed solid case of possible kilonova emission of GRB 211211A. We find the following interesting results.
  • 1.
@@
    Kilonova-like light curves are calculated for three models at different temperatures (4, 5, and 6 GK), and it is found that Model-B shows the brightest emission at each given temperature due to its larger ejecta mass. By comparing with adopting different temperatures of each model, the kilonova-like emission at $T=4$ GK is brighter than that of $T=5$ GK and $T=6$ GK, which is likely due to the higher $\rm{}^{56}Ni$ production at $T=4$ GK.
  • 3.
@@
    By comparing the calculated kilonova-like emission with that of afterglow-subtracted observations of possible kilonova emission in GRB 211211A, the merger of an NS and a WD cannot be ruled out as the origin of GRB 211211A to produce the possible kilonova emission if the remnant of the WD-NS merger is a supramassive or stable magnetar. Otherwise, it is difficult to explain the early possible kilonova emission following GRB 211211A by invoking the merger of a WD and an NS.
@@
We also compare our results with those of other works derived from different models. Zenati et al. (2020) adopted 3D hydrodynamical simulations and 2D hydrodynamical-thermonuclear simulations to model the mergers of NSs with CO WDs, which produce faint transient events. In their results, the predicted transient events are much fainter and faster evolving than those of typical Type Ia SNe. For the peak luminosity, the luminosity at $T=4$ GK in our result is not much different from that of their work, with units of magnitude. However, in terms of the time of the peak, the peak time of our results is earlier than that of their work. The reasons for such a difference may be due to different heating rates or different models for calculating the radiative transfer or differences in expansion velocity and ejecta masses. In Morán-Fraile et al. (2024), the 3D-magnetohydrodynamics (MHD) simulations of the merger were conducted by using the magnetohydrodynamic moving mesh code AREPO. They claimed that the luminosity of the transients produced by NS-WD mergers could be larger than that of typical kilonovae ($\sim 10^{41}\rm~erg\,s^{-1}$), which is also consistent with our results from the aspect of the total luminosity.
@@
In any case, it is difficult to firmly identify the origin of the GRB 211211A-like event only based on its multiwavelength kilonova-like emission. The best way to identify the merger system is to combine with other observational characteristics, such as direct GW detection, or more detailed spectroscopic analysis to look for spectral line features produced by heavier elements, as seen in examples like AT2017gfo, AT2018kzr, and AT2023vfi (Gillanders et al., 2020, 2023; Gillanders & Smartt, 2025; Liu et al., 2025).
@@
Moreover, GRB 230307A is another interesting event that is similar to that of GRB 211211A (Gillanders et al., 2023; Du et al., 2024; Levan et al., 2024; Yang et al., 2024; Sun et al., 2025), and the characteristic of soft X-ray emission suggests that the central engine of GRB 230307A is a magnetar from a binary NS merger (Du et al., 2024; Zhong et al., 2024; Sun et al., 2025). Wang et al. (2024) proposed that the origin of GRB 230307A is also an NS–WD merger, and the late-time kilonova-like emission is powered by spin-down of the magnetar and the radioactive decay of $\rm{}^{56}Ni$. More interestingly, the prominent emission feature of GRB 230307A at $\sim 2.1\mu m$ detected by James Webb Space Telescope is suggested to be likely attributed to Te III with $A=128$ (Gillanders & Smartt, 2025). Such an element is much heavier than the elements obtained by our simulations from the merger of an NS and a WD. If such an emission line in GRB 230307A is indeed real, it means that at least GRB 230307A originated from the merger of an NS and a WD is very difficult to explain the observed emission line of Te III in the spectrum. In other words, once one detects the emission line of heavier elements ($A>90$) in the spectrum of a GRB 230307A-like event in the feature, the origin of the mergers of an NS and a WD would be ruled out.
@@
Meanwhile, it is worth noting that the system of NS-WD merger is not the only progenitor to explain the type of LGRBs associated with possible kilonova emission. A number of other possible origins are also proposed, such as neutron star-neutron star mergers (NS-NS; Gottlieb et al. (2023)), NS-BH mergers (Rastinejad et al., 2022), WD-BH mergers (Fryer et al., 1999; Lloyd-Ronning et al., 2024), massive accretion disks around black hole (Gottlieb et al., 2025), and low-mass collapsar as the progenitor (Zenati et al., 2020; An et al., 2023; Barnes & Metzger, 2023; Dean & Fernández, 2024). Further multimessenger observations will be essential to distinguish between these progenitor scenarios and to establish a more comprehensive understanding of the origin of this type of LGRBs.
[…]